DO 1. KZEP-FM is a San Antonio Radio station that claims to be a Classic Rock station (they arent oldies if remember when
Answer #1

We know, wave length (λ)= c(speed of light)/ ν(frequency of the wave)…………………….(1)

C= 3×108m/s

ν= 106.7 MHz= 106.7×106 Hz (given)

Putting these values in equation (1):

λ= (108)/106.7×106 m

λ= 2.811 m

Hence, wavelength of that radio wave is 2.811 m.